OpenStudy (anonymous):

5x+2(x+6) A)7x+6 B)7x^2+12 C)7x+12 D)7x(x+6)

OpenStudy (anonymous):

7x + 12

OpenStudy (shiree19):

C

OpenStudy (shiree19):

5x + 2 = 7x; 6 * 2 = 12; 7x + 12 = C

OpenStudy (anonymous):

5x stays the same and 2(x + 6) becomes 2x + 12: 5x + 2x + 12 (now simplify it...) Answer: 7x + 12 C.